Mr Lee is now ? years old.his current age is the square of his age twenty years ago .Find the current age of Mr Lee.
Answer by ikleyn(52788)

.
x = (x-20)^2


Answer.  x = 25  (obtained mentally).


Algebra solution

    x = x^2 - 40x + 400


    x^2 -41x + 400 = 0


    (x-25)*(x-16) = 0


    Of two roots of this quadratic equation  x= 25  and  x= 16  only  x= 25  remains positive "20 years ago".